OpenStudy (anonymous):

What are methods for factoring polynomials? (finding real roots, not complex ones) I know some: 1) factor common terms 2) special products 3) factoring by grouping Are there any other ways for factoring?

OpenStudy (anonymous):

and ofcourse, completing the square

OpenStudy (swissgirl):

well you can do long division.

OpenStudy (anonymous):

That is true. But you always need to find a first factor for that

OpenStudy (swissgirl):

truuueeeeeeeee

OpenStudy (swissgirl):

quadratic formula for parabola ?

OpenStudy (swissgirl):

idk cant think of more

OpenStudy (anonymous):

True but that's the same thing as "completing the square"

OpenStudy (anonymous):

I know there are more because my math teacher said so

OpenStudy (swissgirl):

not really. Completing the square in my opinion is a different method

OpenStudy (anonymous):

When you complete the square on a general polynomial ax^2 + bx + c, you get the quadratic formula

OpenStudy (cwrw238):

you can factor a quadratic by inspecting the factors of numbers involved eg x^2 - 4x + 3 factors to (x +/- a)(x +/- b) ab must = + 3 and a + b must = -4 so a = -1 and b = -3 and factors are (x - 1)(x - 3)

OpenStudy (anonymous):

What is that method called? I know it but don't know it's name. Are there more methods?

OpenStudy (anonymous):

I want to know all :p

OpenStudy (cwrw238):

for larger polynomials there the factor theorem inspection of the coefficients of x^3 and last coefficient can help you make a good guess eg x^3 + 5x^2 - 2x - 15 factors are likely to be x +- 1, +/-3, +/-5 or +/- 15 check these out by substitution in the polynomial when you have found one use long or synthetic division to find other factors

OpenStudy (anonymous):

I know the factor theorem. But why are those factors likely?

OpenStudy (cwrw238):

can't remember the name of the above method of likely factors - there is a nme for it. that are likely because the last term is 15

OpenStudy (cwrw238):

3*5 = 15 , 1 *15 = 15

OpenStudy (cwrw238):

oh yes - its the rational root theorem

OpenStudy (cwrw238):

you are familiar with difference of 2 squares i expect?

OpenStudy (cwrw238):

thats a special case i guess

OpenStudy (anonymous):

yes, a binomial multiplied by it's conjugate, right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

yeah, but that falls under "special product"

OpenStudy (cwrw238):

i don't know if there is a name for the factoring of quadratics (trinomials) into 2 binomials.

OpenStudy (anonymous):

How do computers do it? What is the algorithm?

OpenStudy (anonymous):

(I'm a coder) :p

OpenStudy (cwrw238):

no idea

OpenStudy (anonymous):

But can you solve any polynomial with those methods? And how do you know if they are irreducible?

OpenStudy (anonymous):

factor, not solve, sorry

hartnn (hartnn):

u can factor by finding roots , like if 2 is a root then (x-2) is a factor.

OpenStudy (anonymous):

Yes, but that's the factor theorem

OpenStudy (anonymous):

and how do you solve an equation like ax^16 + bx^15+cx^14+... = 0?

OpenStudy (anonymous):

Rational Root theorem

OpenStudy (anonymous):

Descartes Rule of Signs

OpenStudy (anonymous):

Wolfram

OpenStudy (anonymous):

Root finding algorithms